# Privacy in Speech Technology

### v3

@@ VIII Legal and societal landscape
While we have not (yet) seen breaches related to speech technology with consequences of comparable magnitude, these parallels highlight the potential effect of breaches. Scandals directly related to speech technology include eavesdropping on private persons by employees and contractors of service providers [5, 6, 221].
@@
Section VII-C also makes the point that service and technology providers have clear short-term incentives that conflict with users’ preferences for privacy. We can easily find examples where users do not have real choices in protecting their privacy. For example, suppose all friends and family of a user use a particular platform for social and voice interaction even if its privacy configuration is inadequate [222]. The individual user faces then the choice of either disconnecting from their social network or compromising privacy. This applies to all users with a sufficiently large portion of their communities participating in that platform. This is a variation of the prisoner’s dilemma, where no single user has sufficient incentive to transition to a solution that would be best for everyone.
@@
On the individual level, the examples of potential exploits in table II cause clear damage to individuals, including psychological harm in particular (e.g. [90]). Additional dangers include stalkers [223, 224, 225] and oppression [222]. While such individual damage is “small” on the societal level, their prevalence makes their joint impact significant [8].
@@
These examples demonstrate the inherent need for society-level regulation of speech technology with respect to privacy. Governments have already responded to this need, with the European Union spearheading the process with the General Data Protection Regulation (GDPR) [226, 227], with the State of California following soon thereafter with the California Consumer Privacy Act (CCPA) of 2018 [228]. While these laws cover only a small percentage of the global population, as cloud services typically operate globally, they need the capability to follow local laws. In many cases, it can be easiest to apply the strictest laws on all users, such that the most strict laws benefit the privacy of all users. Service providers have, therefore, widely adopted the requirements of GDPR and CCPA, and that has likely had a large impact also on users outside the scope of these regulations. More recently, other countries have followed suit, for example through the “Personal Information Protection Law of the People’s Republic of China” [229] and “Lei Geral de Proteção de Dados” (Brazilian General Data Protection Law, LGPD) [230].
@@
With respect to regulation, an important consequence of the objective measures of privacy in section VI is that our tools and measurements will give as an output only statistical characterizations of the accuracy of estimated attributes, but they can never give absolute confidence. This is in stark contrast with the concept of unique identifiability used in legal documents, such as the General Data Protection Regulation (GDPR) by the European Union [226], which does not explicitly leave room for statistical uncertainty. This is reflected for example in the Guidelines for virtual voice assistants by the European Data Protection Board, which states that: [89, page 13, §31]
@@
This leaves the interpretation open. It is possible to argue that it is never possible to obtain absolute confidence in speaker identification such that the GDPR is never triggered. It is also possible to argue that all voice data contains personal information that can be used to uniquely identify a person, such that all processing must have a valid legal basis. Both interpretations lead to absurdity, which suggests that the truth must lie somewhere in the middle. In fact, the GDPR in practice requires (see [89, page 4]) that the design process of voice assistants includes a data protection impact assessment, where the risks and consequences are evaluated such that the designer can take appropriate precautions to preserve privacy. Authors of the GDPR are thus clearly aware that it is impossible to give absolute guarantees of privacy but that the impact assessment (i.e., objective measures of privacy) must necessarily be based on statistical measures, even if such measures have not been defined. Fortunately, first mathematical characterizations of the legal criteria have started to appear [169, 170].
However, even in the absence of legal requirements, the current author would encourage all researchers, developers, data analysts, project managers, and policymakers to perform impact assessments in data projects, data management, and data policies, anytime when there is or could be the slightest concern that the dataset imposes risks to privacy. A practical tool to aid impact assessment is the “Data Ethics Decision Aid (DEDA)” that formalizes and enumerates all the questions that a representative group of stakeholders should review, discuss, and consider [231].
While governments are in the process of regulating privacy, corporations and non-governmental organizations have also realized that proper privacy is an opportunity. For example, the Open Voice Network seeks to develop and standardize open technical standards and ethical guidelines for voice assistance [232] and the MyData Global seeks to help people and organizations to benefit from personal data in a human-centric way [233, 234]. Within the research community, the author of this paper has been involved in establishing a special interest group within the International Speech Communication Association (ISCA) devoted to “Security and Privacy in Speech Communication” [235]. It is as far as we know, the world-wide largest community focused on this topic.
@@ IX Discussion and conclusions
This paper is a tutorial on privacy for speech technology. Its most notable novelty is an exhaustive categorization of threats (see figs. 1 and IV). Protections against those threats are further categorized according to whether they relate to the target information or side information. The pertinent difference is that transmitting target information is the whole purpose of communication, and there is not much we can do to protect it other than encryption. With side-information, that is, all the other information that is bundled into a speech, like health status and gender identity, we have a much larger arsenal of protections. The primary approach is, however, to remove as much of the side information as possible as early as possible. As the target information is all the communication we need, all side-information should be removed to the extent possible. Such removal, however, rapidly demonstrates that paralinguistic information like speech style is often very useful in conveying the intended message. It is thus not always clear what constitutes the legitimate, target information.
The first conclusion from this paper is that the range of possible threats to privacy is vast. Each agent – be it human or device – participating in an interaction as well as the acoustic pathways and network connection through which they are connected, is a potential attack surface. Any actor that can interact with the other agent or listen to the connection is a potential eavesdropper. Since we define privacy as a scenario where an agent is authorized with some access but over-exceeds that authorization (intentionally or inadvertently), we cannot just cut connections but need more refined designs and methodologies. We thus need to dynamically adjust access according to need. Conversely, systems must actively monitor the privacy status to determine appropriate actions.
Second, we find that privacy and ethics are largely overlapping challenges. Our ethical values govern our preferences for privacy. Most potential breaches of ethics in speech technology are based on breaching privacy. That means that we need a society-wide ethical discussion about what is allowed with respect to user privacy. Such discussions are needed to prevent a Cambridge Analytica-style scandal for speech technology [219].
A third implication of this paper is that, while research in this field has started to accelerate only recently, there is already a substantial body of research available. The research is not, however, mature but in a phase of rapid development, and there are important sub-areas that have not yet seen much work. This makes it a fruitful area for research as we can expect important understanding to be discovered in the coming years.
@@
While management of acquiring informed consent has established traditions and best practices for most interface types [236], speech, audio, and ambient systems are notably unique. Namely, acoustic information is a time-varying stream. Reading out a pages-long consent form before an interaction can start is clearly much too obtrusive and unnecessarily detailed. Privacy requirements also vary over time. Consent should thus be acquired on an actual need basis. In addition to being more usable, it would also make choices better connected to the actual needs since consent is acquired only once it is needed.
@@
The available theoretical metrics reflect privacy with respect to a finite dataset, whereas speech is an open-ended stream of data. The consequence is that, in theory, we can resolve any private attribute or identity, provided that it has a unique probability distribution and we have a sufficiently long observation. We would thus need methodologies for characterizing the effect of the length of observation on privacy.
@@
The metrics discussed in section VI are all related to specific categories of private information and in particular, we can provide protection only to identified threats. For example, we can measure the threat to privacy related to health information, but that does not say anything about the threat related to information about ethnic background. We thus need methods for evaluating privacy jointly with respect to all categories of private information except the target information.
@@
The biometrics community has a long tradition of characterizing attacks as, for example, opportunistic and motivated attacks, as well as in terms of the potential rewards of such attacks [237]. Characterization of speech-specific attacks has received much less attention, though figs. 1 and IV in this paper are attempts in that direction. Further work is needed to gain privacy-protections on a systems level.
@@
Privacy research is categorically focused on personal and user-centric privacy. However, speech is, by definition, communication between multiple agents, and when exposed, threatens all participants simultaneously. This is not an issue from a legal point of view, because privacy protections apply to all individual users equally. However, from an authorization and consent management perspective, this is an underappreciated issue. If user A records a discussion with user B, then both clearly have some level of ownership and privacy requirements on that recording. Another case is smart technology with multiple users, like smart TVs; even if one user has consented to data collection, that does not mean that others would agree [222]. We do not yet have any widely accepted standard approaches for handling privacy, ownership, and consent in such multi-user scenarios.
@@
If we could disentangle all categories of speech information as in fig. 14, it would be easy to anonymize each category to an appropriate degree. This approach thus seemingly solves all our problems. The issue is that we do not yet have sufficiently sophisticated methods for such disentanglement. The difficulty in developing disentanglement algorithms is that information categories in table I are vaguely and heuristically defined and there is significant overlap between them. We cannot even demonstrate that this would be a complete list of information categories. Without exact definitions of those categories, we have no hope of developing methods for them. An alternative approach is to use representation learning methods to create unsupervised clustering of information categories, thus circumventing the need to define categories explicitly. The compromise is that we cannot guarantee that the learned representations correspond to heuristically meaningful categories. Still, since disentanglement is the ideal solution, that should continue to be a central focus of research.
@@
In conclusion, threats and breaches of privacy have significant negative consequences on individual, societal, ethical, and economic levels. While further improvements in smart technology are expected to improve the utility of the technology, it will likely also introduce new threats. The protection of privacy in speech technology has thus been important already for a long time, and its importance is increasing. Fortunately, research in the area has picked up speed, and this tutorial presents the most important concepts, approaches, and methodology. It is, however, likely that fundamental results and new technologies will be introduced in the near future. This is thus an exciting time for researchers in the area.

### v4

@@ VIII Legal and societal landscape
While we have not (yet) seen breaches related to speech technology with consequences of comparable magnitude, these parallels highlight the potential effect of breaches. Scandals directly related to speech technology include eavesdropping on private persons by employees and contractors of service providers [Lynskey2019GuardianPrivacyScandal, hern2019, day2019].
@@
Section VII-C also makes the point that service and technology providers have clear short-term incentives that conflict with users’ preferences for privacy. We can easily find examples where users do not have real choices in protecting their privacy. For example, suppose all friends and family of a user use a particular platform for social and voice interaction, even if its privacy configuration is inadequate [backstrom2025beyonduser]. The individual user then faces the choice of either disconnecting from their social network or compromising privacy. This applies to all users with a sufficiently large portion of their communities participating in that platform. This is a variation of the prisoner’s dilemma, where no single user has sufficient incentive to transition to a solution that would be best for everyone.
@@
On the individual level, the examples of potential exploits in table II cause clear damage to individuals, including psychological harm in particular (e.g. [natarajan2021prioris]). Additional dangers include stalkers [lu2019abusers, hulsen2022stalkers, avast2021stalkerware] and oppression [backstrom2025beyonduser]. While such individual damage is “small” on the societal level, their prevalence makes their joint impact significant [levy2020intimate].
@@
These examples demonstrate the inherent need for society-level regulation of speech technology with respect to privacy. Governments have already responded to this need, with the European Union spearheading the process with the General Data Protection Regulation (GDPR) [eu2016gdpr, nautsch19c_interspeech], with the State of California following soon thereafter with the California Consumer Privacy Act (CCPA) of 2018 [california2018ccpa]. While these laws cover only a small percentage of the global population, as cloud services typically operate globally, they need the capability to follow local laws. In many cases, it can be easiest to apply the strictest laws on all users, such that the most strict laws benefit the privacy of all users. Service providers have, therefore, widely adopted the requirements of GDPR and CCPA, and that has likely had a significant impact also on users outside the scope of these regulations. More recently, other countries have followed suit, for example, through the “Personal Information Protection Law of the People’s Republic of China” [china2021pip] and “Lei Geral de Proteção de Dados” (Brazilian General Data Protection Law, LGPD) [brazil2020lgpd].
@@
With respect to regulation, an important consequence of the objective measures of privacy in section VI is that our tools and measurements will give as an output only statistical characterizations of the accuracy of estimated attributes, but they can never give absolute confidence. This is in stark contrast with the concept of unique identifiability used in legal documents, such as the General Data Protection Regulation (GDPR) by the European Union [eu2016gdpr], which does not explicitly leave room for statistical uncertainty. This is reflected, for example, in the Guidelines for virtual voice assistants by the European Data Protection Board, which states that: [edpb2021guidelines, page 13, §31]
@@
This leaves the interpretation open. It is possible to argue that it is never possible to obtain absolute confidence in speaker identification such that the GDPR is never triggered. It is also possible to argue that all voice data contains personal information that can be used to uniquely identify a person, such that all processing must have a valid legal basis. Both interpretations lead to absurdity, which suggests that the truth must lie somewhere in the middle. In fact, the GDPR in practice requires (see [edpb2021guidelines, page 4]) that the design process of voice assistants includes a data protection impact assessment, where the risks and consequences are evaluated such that the designer can take appropriate precautions to preserve privacy. Authors of the GDPR are thus clearly aware that it is impossible to give absolute guarantees of privacy but that the impact assessment (i.e., objective measures of privacy) must necessarily be based on statistical measures, even if such measures have not been defined. Fortunately, the first mathematical characterizations of the legal criteria have started to appear [vauquier25_interspeech, cohen2020towards].
However, even in the absence of legal requirements, the current author would encourage all researchers, developers, data analysts, project managers, and policymakers to perform impact assessments in data projects, data management, and data policies, anytime when there is or could be the slightest concern that the dataset imposes risks to privacy. A practical tool to aid impact assessment is the “Data Ethics Decision Aid (DEDA)” that formalizes and enumerates all the questions that a representative group of stakeholders should review, discuss, and consider [franzke2021data].
While governments are in the process of regulating privacy, corporations and non-governmental organizations have also realized that proper privacy is an opportunity. For example, the Open Voice Network seeks to develop and standardize open technical standards and ethical guidelines for voice assistance [ovn2023] and the MyData Global seeks to help people and organizations to benefit from personal data in a human-centric way [mydata2023, poikola2020mydata]. Within the research community, the author of this paper has been involved in establishing a special interest group within the International Speech Communication Association (ISCA) devoted to “Security and Privacy in Speech Communication” [spsc2023]. It is, as far as we know, the world’s largest community focused on this topic.
@@ IX Discussion and conclusions
This paper is a tutorial on privacy for speech technology. Its most notable novelty is an exhaustive categorization of threats (see figs. 1 and IV). Protections against those threats are further categorized according to whether they relate to the target information or side information. The pertinent difference is that transmitting target information is the whole purpose of communication, and there is not much we can do to protect it other than encryption. With side information, that is, all the other information that is bundled into a speech, like health status and gender identity, we have a much larger arsenal of protections. The primary approach is, however, to remove as much of the side information as possible as early as possible. As the target information is all the communication we need, all side information should be removed to the extent possible. Such removal, however, rapidly demonstrates that paralinguistic information like speech style is often instrumental in conveying the intended message. It is thus not always clear what constitutes the legitimate, target information.
The first conclusion from this paper is that the range of possible privacy threats is vast. Each agent – be it human or device – participating in an interaction, as well as the acoustic pathways and network connection through which they are connected, is a potential attack surface. Any actor that can interact with the other agent or listen to the connection is a potential eavesdropper. Since we define privacy as a scenario where an agent is authorized with some access but over-exceeds that authorization (intentionally or inadvertently), we cannot just cut connections but need more refined designs and methodologies. We thus need to dynamically adjust access according to need. Conversely, systems must actively monitor the privacy status to determine appropriate actions.
Second, we find that privacy and ethics are largely overlapping challenges. Our ethical values govern our preferences for privacy. Most potential breaches of ethics in speech technology are based on breaching privacy. That means that we need a society-wide ethical discussion about what is allowed with respect to user privacy. Such discussions are necessary to prevent a Cambridge Analytica-style scandal for speech technology [heikkila2023cambridge_ai].
A third implication of this paper is that, while research in this field has started to accelerate only recently, there is already a substantial body of research available. The research is not, however, mature but in a phase of rapid development, and there are important sub-areas that have not yet seen much work. This makes it a fruitful area for research, as we can expect important understanding to be discovered in the coming years.
@@
While management of acquiring informed consent has established traditions and best practices for most interface types [friedman2000informed], speech, audio, and ambient systems are notably unique. Namely, acoustic information is a time-varying stream. Reading out a pages-long consent form before an interaction can start is clearly much too obtrusive and unnecessarily detailed. Privacy requirements also vary over time. Consent should thus be acquired on an actual need basis. In addition to being more usable, it would also make choices better connected to the actual needs, since consent is acquired only once it is needed.
@@
The available theoretical metrics reflect privacy with respect to a finite dataset, whereas speech is an open-ended stream of data. The consequence is that, in theory, we can resolve any private attribute or identity, provided that it has a unique probability distribution and we have a sufficiently long observation. We would thus need methodologies for characterizing the effect of the length of observation on privacy. The current author has recently made first attempts in this direction [backstrom2025privacydisclosure].
@@
The metrics discussed in section VI are all related to specific categories of private information, and in particular, we can provide protection only to identified threats. For example, we can measure the threat to privacy related to health information, but that does not say anything about the threat related to information about ethnic background. We thus need methods for evaluating privacy jointly with respect to all categories of private information except the target information.
@@
The biometrics community has a long tradition of characterizing attacks as, for example, opportunistic and motivated attacks, as well as in terms of the potential rewards of such attacks [roberts2007biometric]. Characterization of speech-specific attacks has received much less attention, though figs. 1 and IV in this paper are attempts in that direction. Further work is needed to gain privacy protections on a system-level.
@@
Privacy research is categorically focused on personal and user-centric privacy. However, speech is, by definition, communication between multiple agents, and when exposed, threatens all participants simultaneously. This is not an issue from a legal point of view, because privacy protections apply to all individual users equally. However, from an authorization and consent management perspective, this is an underappreciated issue. If user A records a discussion with user B, then both clearly have some level of ownership and privacy requirements on that recording. Another case is smart technology with multiple users, like smart TVs; even if one user has consented to data collection, that does not mean that others would agree [backstrom2025beyonduser]. We do not yet have any widely accepted standard approaches for handling privacy, ownership, and consent in such multi-user scenarios.
@@
If we could disentangle all categories of speech information as in fig. 14, it would be easy to anonymize each category to an appropriate degree. This approach thus seemingly solves all our problems. The issue is that we do not yet have sufficiently sophisticated methods for such disentanglement. The difficulty in developing disentanglement algorithms is that information categories in table I are vaguely and heuristically defined, and there is significant overlap between them. We cannot even demonstrate that this would be a complete list of information categories. Without exact definitions of those categories, we have no hope of developing methods for them. An alternative approach is to use representation learning methods to create unsupervised clustering of information categories, thus circumventing the need to define categories explicitly. The compromise is that we cannot guarantee that the learned representations correspond to heuristically meaningful categories. Still, since disentanglement is the ideal solution, that should continue to be a central focus of research.
@@
In conclusion, threats and breaches of privacy have significant negative consequences on individual, societal, ethical, and economic levels. While further improvements in smart technology are expected to improve the utility of the technology, it will likely also introduce new threats. The protection of privacy in speech technology has thus been important already for a long time, and its importance is increasing. Fortunately, research in the area has picked up speed, and this tutorial presents the most important concepts, approaches, and methodologies. It is, however, likely that fundamental results and new technologies will be introduced in the near future. This is thus an exciting time for researchers in the area.
@@ Acknowledgments
Feedback from the anonymous reviewers was outstanding in both quality and extent. It pains me that I can never adequately acknowledge their exceptional work.
@@
[Uncaptioned image] Tom Bäckström D.Sc. (tech.) is an associate professor at Aalto University, Finland (2019-). He obtained his Master’s and Doctoral degrees at Helsinki University of Technology (the predecessor of Aalto) in 2001 and 2004, respectively. During his time at International Audio Laboratories Erlangen, Germany (2008-2016), he made contributions to several international speech and audio coding standards such as MPEG USAC and 3GPP EVS, and became a professor (W2) at Friedrich-Alexander University Erlangen-Nürnberg (FAU) (2012-2016). Before his current position, he was a professor practice at Aalto University (2016-2019). He is the President (2025–2027) and Board Member (2023–2027) of the International Speech Communication Association (ISCA). He was the initiator and chair of ISCA SIG “Security and Privacy in Speech Communication” (2019-2022). His current research interests include privacy, trust, coding, enhancement, and transmission of speech as well as machine learning.